# The quadratic linking degree

### v5

@@ Appendix A An explicit definition of the residue morphisms of Milnor-Witt $K$-theory
Let $m\in\mathbb{Z}$ and $L$ be an $F$-vector space of dimension $1$. The $L$-twisted $m$-th Milnor-Witt $K$-theory abelian group of $F$, denoted $K^{\operatorname{MW}}_{m}(F,L)$, is the tensor product of the $\mathbb{Z}[F^{*}]$-modules $K^{\operatorname{MW}}_{m}(F)$ and $\mathbb{Z}[L\setminus\{0\}]$ (the scalar product of $K^{\operatorname{MW}}_{m}(F)$ being $(\sum_{f\in F^{*}}n_{f}\lambda_{f})\cdot\alpha=\sum_{f\in F^{*}}n_{f}\langle f\rangle\alpha$).
@@
$+\eta\chi^{\operatorname{odd}}(m_{1}m_{2})[-1,\overline{u_{1}}]+\eta\chi^{\operatorname{odd}}(m_{1})[-1,\overline{u_{2}}]$
@@ B.1. Definitions and first properties
Note that by definition, for all $i\in\mathbb{N}_{0}$ and $j\in\mathbb{Z}$, $\mathcal{C}^{i}(\operatorname{Spec}(F),\underline{K}^{\operatorname{MW}}_{j})=K^{\operatorname{MW}}_{j}(F)$ if $i=0$, to 00 otherwise, hence $H^{i}(\operatorname{Spec}(F),\underline{K}^{\operatorname{MW}}_{j})=K^{\operatorname{MW}}_{j}(F)$ if $i=0$, to 00 otherwise.
@@
Note that it follows from this theorem that for all $n,i\in\mathbb{N}_{0}$ and $j\in\mathbb{Z}$, $H^{i}(\mathbb{A}^{n}_{F},\underline{K}^{\operatorname{MW}}_{j})$ is canonically isomorphic to $H^{i}(\operatorname{Spec}(F),\underline{K}^{\operatorname{MW}}_{j})$ hence to $K^{\operatorname{MW}}_{j}(F)$ if $i=0$, to 00 otherwise.
@@
A boundary triple is a $5$-tuple $(Z,i,X,j,U)$, or abusively a triple $(Z,X,U)$, with $i:Z\to X$ a closed immersion and $j:U\to X$ an open immersion such that the image of $U$ by $j$ is the complement in $X$ of the image of $Z$ by $i$, where $Z,X,U$ are smooth $F$-schemes of pure dimensions. We denote by $d_{Z}$ and $d_{X}$ the dimensions of $Z$ and $X$ respectively and by $\nu_{Z}$ the determinant of the normal sheaf of $Z$ in $X$.
@@ B.2. The Rost-Schmid groups of punctured affine spaces
Note that elements of $\mathcal{C}^{i}(X,\underline{K}^{\operatorname{MW}}_{j}\{\mathcal{L}\otimes\mathcal{L}\})$ are of the form $\sum_{x\in I}m_{x}\otimes t_{x}$ with $I$ a finite subset of $X^{(i)}$, $m_{x}\in K^{\operatorname{MW}}_{j-i}(\kappa(x))$ and $t_{x}\in\mathbb{Z}[(\nu_{x}\otimes(\mathcal{L}\otimes\mathcal{L})_{|x})\setminus\{0\}]$. Let $x\in I$. Since $\nu_{x}\otimes(\mathcal{L}\otimes\mathcal{L})_{|x}$ is a $\kappa(x)$-vector space of dimension $1$, there exist $n_{x}\in K^{\operatorname{MW}}_{j-i}(\kappa(x))$ and $s_{x}\in(\nu_{x}\otimes(\mathcal{L}\otimes\mathcal{L})_{|x})\setminus\{0\}$ such that $m_{x}\otimes t_{x}=n_{x}\otimes s_{x}$. By definition of $K^{\operatorname{MW}}_{j-i}(\kappa(x),\nu_{x})$, there exist $h_{x}\in K^{\operatorname{MW}}_{j-i}(\kappa(x),\nu_{x})$ and $l_{x},r_{x}\in\mathcal{L}_{|x}\setminus\{0\}$ such that $n_{x}\otimes s_{x}=h_{x}\otimes(l_{x}\otimes r_{x})$. Since $\mathcal{L}_{|x}$ is a $\kappa(x)$-vector space of dimension $1$, there exists $v_{x}\in\kappa(x)^{*}$ such that $r_{x}=v_{x}l_{x}$. It follows that $h_{x}\otimes(l_{x}\otimes r_{x})=\langle v_{x}\rangle h_{x}\otimes(l_{x}\otimes l_{x})$. Denoting $k_{x}:=\langle v_{x}\rangle h_{x}$, we get that elements of $\mathcal{C}^{i}(X,\underline{K}^{\operatorname{MW}}_{j}\{\mathcal{L}\otimes\mathcal{L}\})$ are of the form $\sum_{x\in I}k_{x}\otimes(l_{x}\otimes l_{x})$ with $I$ a finite subset of $X^{(i)}$, $k_{x}\in K^{\operatorname{MW}}_{j-i}(\kappa(x),\nu_{x})$ and $l_{x}\in\mathcal{L}_{|x}\setminus\{0\}$.
@@
induced by the isomorphism $\det(\mathcal{N}_{\{0\}/\mathbb{A}^{2}_{F}})\to\mathcal{O}_{\{0\}}\otimes\mathcal{O}_{\{0\}}$ which sends $\overline{u}^{*}\wedge\overline{v}^{*}$ to $1\otimes 1$, where $\mathbb{A}^{2}_{F}=\operatorname{Spec}(F[u,v])$ (see Definition B.15) and of the canonical isomorphism (which sends $\eta^{2}$ to $1$)
@@
induced by the isomorphism $\det(\mathcal{N}_{\{0\}/\mathbb{A}^{4}_{F}})\to\mathcal{O}_{\{0\}}\otimes\mathcal{O}_{\{0\}}$ which sends $\overline{x}^{*}\wedge\overline{y}^{*}\wedge\overline{z}^{*}\wedge\overline{t}^{*}$ to $1\otimes 1$, where $\mathbb{A}^{4}_{F}=\operatorname{Spec}(F[x,y,z,t])$ (see Definition B.15) and of the canonical isomorphism (which sends $\eta^{2}$ to $1$)
@@ B.3. The intersection product of oriented divisors
Let $c_{1},c_{2}$ be oriented divisors in $X$, i.e. $c_{1},c_{2}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$. The intersection product of $c_{1}$ with $c_{2}$, denoted by $c_{1}\cdot c_{2}$, is $\langle-1\rangle$-commutative:
@@
Let $X$ be a smooth $F$-scheme. Let $D_{1},D_{2}$ be smooth integral divisors in $X$ such that $D_{1}\cap D_{2}$ is of codimension $2$ in $X$. For all $i\in\{1,2\}$, let $g_{i}$ be a local parameter for $D_{i}$, i.e. $g_{i}$ is a uniformizing parameter for $\mathcal{O}_{X,D_{i}}$. The intersection product of $1\otimes\overline{g_{1}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ (over the generic point of $D_{1}$) with $1\otimes\overline{g_{2}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ (over the generic point of $D_{2}$) is the class in $H^{2}(X,\underline{K}^{\operatorname{MW}}_{2})$ of the sum over the generic points $x$ of the irreducible components of $D_{1}\cap D_{2}$ of $(m_{x})_{\epsilon}\langle u_{x}\rangle\otimes(\overline{\pi_{x}}^{*}\otimes\overline{g_{1}}^{*})$ (over the point $x$), where $\pi_{x}$ is a uniformizing parameter for $\mathcal{O}_{X,x}/(g_{1})$, $u_{x}$ is a unit in $\mathcal{O}_{X,x}/(g_{1})$ and $m_{x}\in\mathbb{Z}$ such that $g_{2}=u_{x}\pi_{x}^{m_{x}}\in\mathcal{O}_{X,x}/(g_{1})$.
@@
  • Denoting by $i_{1}:D_{1}\to X$ the inclusion and by $\Theta_{1}:H^{0}(D_{1},\underline{K}^{\operatorname{MW}}_{0})\to H^{0}(D_{1},\underline{K}^{\operatorname{MW}}_{0}\{\nu_{D_{1}}\})$ (where $\nu_{D_{1}}$ is the determinant of the normal sheaf of $D_{1}$ in $X$) the isomorphism which sends $1$ to $1\otimes\overline{g_{1}}^{*}$, check that $1\otimes\overline{g_{1}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ is equal to $(i_{1})_{*}(\Theta_{1}(1))$.
@@
Let $X$ be a smooth $F$-scheme. Let $D_{1},D_{2}$ be smooth integral divisors in $X$ such that $D_{1}\cap D_{2}$ is of codimension $2$ in $X$. For all $i\in\{1,2\}$, let $g_{i}$ be a local parameter for $D_{i}$ and $f_{i}$ be a unit in $\kappa(D_{i})=\mathcal{O}_{X,D_{i}}/\mathfrak{m}_{X,D_{i}}$ such that for all generic points $x$ of irreducible components of $D_{1}\cap D_{2}$, $f_{i}\in\kappa(x)=\mathcal{O}_{X,x}/\mathfrak{m}_{X,x}$ is a unit. The intersection product of $\langle f_{1}\rangle\otimes\overline{g_{1}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ (over the generic point of $D_{1}$) with $\langle f_{2}\rangle\otimes\overline{g_{2}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ (over the generic point of $D_{2}$) is the class in $H^{2}(X,\underline{K}^{\operatorname{MW}}_{2})$ of the sum over the generic points $x$ of the irreducible components of $D_{1}\cap D_{2}$ of $(m_{x})_{\epsilon}\langle f_{1}f_{2}u_{x}\rangle\otimes(\overline{\pi_{x}}^{*}\otimes\overline{g_{1}}^{*})$ (over the point $x$), where $\pi_{x}$ is a uniformizing parameter for $\mathcal{O}_{X,x}/(g_{1})$, $u_{x}$ is a unit in $\mathcal{O}_{X,x}/(g_{1})$ and $m_{x}\in\mathbb{Z}$ such that $g_{2}=u_{x}\pi_{x}^{m_{x}}\in\mathcal{O}_{X,x}/(g_{1})$.

### v6

@@ Appendix A An explicit definition of the residue morphisms of Milnor-Witt $K$-theory
Let $m\in\mathbb{Z}$ and $L$ be an $F$-vector space of dimension 111. The $L$-twisted $m$-th Milnor-Witt $K$-theory abelian group of $F$, denoted $K^{\operatorname{MW}}_{m}(F,L)$, is the tensor product of the $\mathbb{Z}[F^{*}]$-modules $K^{\operatorname{MW}}_{m}(F)$ and $\mathbb{Z}[L\setminus\{0\}]$ (the scalar product of $K^{\operatorname{MW}}_{m}(F)$ being $(\sum_{f\in F^{*}}n_{f}\lambda_{f})\cdot\alpha=\sum_{f\in F^{*}}n_{f}\langle f\rangle\alpha$).
@@
$+\eta\chi^{\operatorname{odd}}(m_{1}m_{2})[-1,\overline{u_{1}}]+\eta\chi^{\operatorname{odd}}(m_{1}m_{2})[-1,\overline{u_{2}}]$
@@ B.1. Definitions and first properties
Note that by definition, for all $i\in\mathbb{N}_{0}$ and $j\in\mathbb{Z}$, $\mathcal{C}^{i}(\operatorname{Spec}(F),\underline{K}^{\operatorname{MW}}_{j})=K^{\operatorname{MW}}_{j}(F)$ if $i=0$, to 0 otherwise, hence $H^{i}(\operatorname{Spec}(F),\underline{K}^{\operatorname{MW}}_{j})=K^{\operatorname{MW}}_{j}(F)$ if $i=0$, to 0 otherwise.
@@
Note that it follows from this theorem that for all $n,i\in\mathbb{N}_{0}$ and $j\in\mathbb{Z}$, $H^{i}(\mathbb{A}^{n}_{F},\underline{K}^{\operatorname{MW}}_{j})$ is canonically isomorphic to $H^{i}(\operatorname{Spec}(F),\underline{K}^{\operatorname{MW}}_{j})$ hence to $K^{\operatorname{MW}}_{j}(F)$ if $i=0$, to 0 otherwise.
@@
[…]
@@ B.2. The Rost-Schmid groups of punctured affine spaces
Note that elements of $\mathcal{C}^{i}(X,\underline{K}^{\operatorname{MW}}_{j}\{\mathcal{L}\otimes\mathcal{L}\})$ are of the form $\sum_{x\in I}m_{x}\otimes t_{x}$ with $I$ a finite subset of $X^{(i)}$, $m_{x}\in K^{\operatorname{MW}}_{j-i}(\kappa(x))$ and $t_{x}\in\mathbb{Z}[(\nu_{x}\otimes(\mathcal{L}\otimes\mathcal{L})_{|x})\setminus\{0\}]$. Let $x\in I$. Since $\nu_{x}\otimes(\mathcal{L}\otimes\mathcal{L})_{|x}$ is a $\kappa(x)$-vector space of dimension 111, there exist $n_{x}\in K^{\operatorname{MW}}_{j-i}(\kappa(x))$ and $s_{x}\in(\nu_{x}\otimes(\mathcal{L}\otimes\mathcal{L})_{|x})\setminus\{0\}$ such that $m_{x}\otimes t_{x}=n_{x}\otimes s_{x}$. By definition of $K^{\operatorname{MW}}_{j-i}(\kappa(x),\nu_{x})$, there exist $h_{x}\in K^{\operatorname{MW}}_{j-i}(\kappa(x),\nu_{x})$ and $l_{x},r_{x}\in\mathcal{L}_{|x}\setminus\{0\}$ such that $n_{x}\otimes s_{x}=h_{x}\otimes(l_{x}\otimes r_{x})$. Since $\mathcal{L}_{|x}$ is a $\kappa(x)$-vector space of dimension 111, there exists $v_{x}\in\kappa(x)^{*}$ such that $r_{x}=v_{x}l_{x}$. It follows that $h_{x}\otimes(l_{x}\otimes r_{x})=\langle v_{x}\rangle h_{x}\otimes(l_{x}\otimes l_{x})$. Denoting $k_{x}:=\langle v_{x}\rangle h_{x}$, we get that elements of $\mathcal{C}^{i}(X,\underline{K}^{\operatorname{MW}}_{j}\{\mathcal{L}\otimes\mathcal{L}\})$ are of the form $\sum_{x\in I}k_{x}\otimes(l_{x}\otimes l_{x})$ with $I$ a finite subset of $X^{(i)}$, $k_{x}\in K^{\operatorname{MW}}_{j-i}(\kappa(x),\nu_{x})$ and $l_{x}\in\mathcal{L}_{|x}\setminus\{0\}$.
@@
induced by the isomorphism $\det(\mathcal{N}_{\{0\}/\mathbb{A}^{2}_{F}})\to\mathcal{O}_{\{0\}}\otimes\mathcal{O}_{\{0\}}$ which sends $\overline{u}^{*}\wedge\overline{v}^{*}$ to $\otimes 1$ 1, where $\mathbb{A}^{2}_{F}=\operatorname{Spec}(F[u,v])$ (see Definition B.15) and of the canonical isomorphism (which sends $\eta^{2}$ to 111)
@@
induced by the isomorphism $\det(\mathcal{N}_{\{0\}/\mathbb{A}^{4}_{F}})\to\mathcal{O}_{\{0\}}\otimes\mathcal{O}_{\{0\}}$ which sends $\overline{x}^{*}\wedge\overline{y}^{*}\wedge\overline{z}^{*}\wedge\overline{t}^{*}$ to $\otimes 1$ 1, where $\mathbb{A}^{4}_{F}=\operatorname{Spec}(F[x,y,z,t])$ (see Definition B.15) and of the canonical isomorphism (which sends $\eta^{2}$ to 111)
@@ B.3. The intersection product of oriented divisors
Let $c_{1},c_{2}$ be oriented divisors in $X$, i.e. $c_{1},c_{2}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$. The intersection product of $c_{1}$ with $c_{2}$, denoted by $c_{1}\cdot c_{2}$, is 1$\langle$-1$\rangle$ - 1 ⟩-commutative:
@@
Let $X$ be a smooth $F$-scheme. Let $D_{1},D_{2}$ be smooth integral divisors in $X$ such that $D_{1}\cap D_{2}$ is of codimension 222 in $X$. For all $i\in\{1,2\}$, let $g_{i}$ be a local parameter for $D_{i}$, i.e. $g_{i}$ is a uniformizing parameter for $\mathcal{O}_{X,D_{i}}$. The intersection product of $1\otimes\overline{g_{1}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ (over the generic point of $D_{1}$) with $1\otimes\overline{g_{2}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ (over the generic point of $D_{2}$) is the class in $H^{2}(X,\underline{K}^{\operatorname{MW}}_{2})$ of the sum over the generic points $x$ of the irreducible components of $D_{1}\cap D_{2}$ of $(m_{x})_{\epsilon}\langle u_{x}\rangle\otimes(\overline{\pi_{x}}^{*}\otimes\overline{g_{1}}^{*})$ (over the point $x$), where $\pi_{x}$ is a uniformizing parameter for $\mathcal{O}_{X,x}/(g_{1})$, $u_{x}$ is a unit in $\mathcal{O}_{X,x}/(g_{1})$ and $m_{x}\in\mathbb{Z}$ such that $g_{2}=u_{x}\pi_{x}^{m_{x}}\in\mathcal{O}_{X,x}/(g_{1})$.
@@
  • Denoting by $i_{1}:D_{1}\to X$ the inclusion and by $\Theta_{1}:H^{0}(D_{1},\underline{K}^{\operatorname{MW}}_{0})\to H^{0}(D_{1},\underline{K}^{\operatorname{MW}}_{0}\{\nu_{D_{1}}\})$ (where $\nu_{D_{1}}$ is the determinant of the normal sheaf of $D_{1}$ in $X$) the isomorphism which sends 111 to $1\otimes\overline{g_{1}}^{*}$, check that $1\otimes\overline{g_{1}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ is equal to $(i_{1})_{*}(\Theta_{1}(1))$.
@@
Let $X$ be a smooth $F$-scheme. Let $D_{1},D_{2}$ be smooth integral divisors in $X$ such that $D_{1}\cap D_{2}$ is of codimension 222 in $X$. For all $i\in\{1,2\}$, let $g_{i}$ be a local parameter for $D_{i}$ and $f_{i}$ be a unit in $\kappa(D_{i})=\mathcal{O}_{X,D_{i}}/\mathfrak{m}_{X,D_{i}}$ such that for all generic points $x$ of irreducible components of $D_{1}\cap D_{2}$, $f_{i}\in\kappa(x)=\mathcal{O}_{X,x}/\mathfrak{m}_{X,x}$ is a unit. The intersection product of $\langle f_{1}\rangle\otimes\overline{g_{1}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ (over the generic point of $D_{1}$) with $\langle f_{2}\rangle\otimes\overline{g_{2}}^{*}\in H^{1}(X,\underline{K}^{\operatorname{MW}}_{1})$ (over the generic point of $D_{2}$) is the class in $H^{2}(X,\underline{K}^{\operatorname{MW}}_{2})$ of the sum over the generic points $x$ of the irreducible components of $D_{1}\cap D_{2}$ of $(m_{x})_{\epsilon}\langle f_{1}f_{2}u_{x}\rangle\otimes(\overline{\pi_{x}}^{*}\otimes\overline{g_{1}}^{*})$ (over the point $x$), where $\pi_{x}$ is a uniformizing parameter for $\mathcal{O}_{X,x}/(g_{1})$, $u_{x}$ is a unit in $\mathcal{O}_{X,x}/(g_{1})$ and $m_{x}\in\mathbb{Z}$ such that $g_{2}=u_{x}\pi_{x}^{m_{x}}\in\mathcal{O}_{X,x}/(g_{1})$.